OpenStudy (anonymous):

If someone picks a card from a well-shuffled 52 card deck, what is the probability that the card is not a face card (jack, queen, king)?

OpenStudy (amistre64):

it is 1 - P(face card)

OpenStudy (amistre64):

P(face card) = 12/52 = 3/13 1 - (3/13) = probability its NOT a face card